• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name V. Sulkosky

    Log entry time 09:40:23 on August 27, 2009

    Entry number 286244

    This entry is a followup to: 286175

    keyword=Re: HRS VDC status

    The missing wires 0-15 in V2 and about 10 wires in U2 near 340 are known.
    These problems are left over from the R-HRS top chamber VDC swap back in April.
    As a comparison, you can check this entry from May during E05-102:

    Wire 80 in V1 is a new problem and should be checked. The left HRS VDC
    looks like there is a mapping problem. One could use the VDC database
    from E05-102, which did not have this problem.

    A copy of this log entry has been emailed to: camsonne,paschke